a 136-inch pipe is cut into two pieces. one piece is three times the length of the other. find the lengths of the two pieces
Answer by checkley77(12569)

x+3x=136
4x=136
x=136/4
x=34 inches for the shorter piece.
3*34=102 inches for the longer piece.
Proof:
34+102=136
136=136